OpenStudy (anonymous):

Phyllis invested 12,000, a portion earning a simple interest rate of 4 1/2% per year and the rest earning a rate of 4% per year. After one year the total interest earned on these investments was $525. How much money did she invest at each rate?

OpenStudy (anonymous):

x*9/200+(12000-x)*4/100=525 9x-8x+96000=105000 x=105000-96000

OpenStudy (anonymous):

x=9000 at 41/2% and 3000 at 4%

OpenStudy (anonymous):

oh pellet. i'm doing it right. i just stopped early... thanks